If A=(-3,4),B=(-1,-2),C=(5,6),D=(x,-4) are vertices of a quadrilateral such that area of the triangle ABD=2 area of the triangle ACD, then x=

Grade:11

2 Answers

sahil
6 years ago
see when you will expand determinants for 2 triangle you get 2+6x=-70-2x again 72=-7x and x=-9if correct click approve
sahil
6 years ago
if you will solve determinant you get 6x+2=2(-70-2x)solve to get x=-71/5if answer is correct please give approval